OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x - 3y = 3 5x - 4y = 3 (4 points) A) x can be any number as there are infinitely many solutions to this system B) There is no x value as there is no solution to this system C) -3 D) 3

OpenStudy (anonymous):

That would be d, if you insert 3 in 4(x) - 3y = 3, 4(3) - 3 (let's insert 3 here too) = 0 you would get 12-9 = 3. The other one works too

OpenStudy (anonymous):

Thank you so much!!